Solve for X: 6x-7(x^1/2)- 3=0
Answer by stanbon(75887)

Solve for X: 6x-7(x^1/2)- 3=0
--------------
Rearrange:
6x - 9sqrt(x) + 2sqrt(x) - 3 = 0
Factor to get:
3sqrt(x)(2sqrt(x) - 3)+(2sqrt(x) - 3) = 0
Factor again to get:
(2sqrt(x)-3)(3sqrt(x) + 1) = 0
---------
sqrt(x) = 3/2 or sqrt(x) = -1/3
----------------------
x = 9/4 or x = 1/9
======================
Check the answers to see if either is extraneous.
======================
Cheers,
Stan H.